第37回環論および表現論シンポジウム(2004年) 第二報

日本学術振興会科学研究費補助金基盤研究 (B)(1)(研究代表者：西田憲司 (信州大学理学部)) による上記シンポジウムのプログラムが決定しましたので，お知らせいたします． なお，懇親会参加を予定されていてまだ連絡されていない方は e-mail にて下記会場責任者宛お知らせ下さい．プログラムの追加・変更等，シンポジウムに関する情報については，環論ホームページ

http://fuji.cec.yamanashi.ac.jp/˜ ring/

をご覧下さい．

プログラム責任者 浅芝秀人（大阪市立大学大学院理学研究科）

- - - - - - - - - - - - 記 - - - - - - - - - - - -

期間：２００ 4 年９月３日（金）〜５日（日）

会場：松本市 M ウィング

(http://www.city.matsumoto.nagano.jp/www cbox/html/chukou/chuou/kyochuo01.html)

〒 390-0811 長野県松本市中央１−１８−１

Tel.: 0263-32-1132

懇親会：９月４日（土） 19:00 より

会場責任者：西田憲司（信州大学理学部）（ｋｅｎｉｓｉｄａ＠ math.shinshu-u.ac.jp）

プログラム責任者： 浅芝秀人（大阪市立大学大学院理学研究科）（ａｓａｓｈｉｂａ＠ sci.osaka-cu.ac.jp）

プログラム

９月３日（金）

09:10/09:40 山形邦夫（東京農工大）正ガロア被覆をもつフロベニウス多元環の安定性について

09:50/10:20 若松隆義（埼玉大）一般傾斜加群から導かれる安定同値だが導来同値でない対称多元環

10:30/11:20 Rickard, Jeremy (Univ. of Bristol)

Derived Categories in Representation Theory, I

11:30/12:00 若松隆義，飛田明彦（埼玉大） Tilting modules and Gorenstein property

12:10/12:40 星野光男，加藤義明（筑波大） A construction of Auslander-Gorenstein rings

14:00/14:30 奥山豪（北大理） On the holonomic rank for A-hypergeometric system

14:40/15:10 浅芝秀人（大阪市大理） Domestic canonical algebras and simple Lie algebras

15:20/16:10 Van den Bergh, Michel (Limburgs Universitair Centrum)

Non-commutative Del Pezzo surfaces, I

16:30/17:00 小池寿俊（沖縄工高専） Morita duality and ring extensions

17:10/17:40 丸林英俊; Brungs, H., Osmanagic, E. （鳴門教育大; Alberta Univ.） Non-commutative valuation rings of crossed product algebras

17:50/18:20 藤田尚昌，酒井洋介 Frobenius full matrix algebras and Gorenstein tiled orders

９月４日（土）

09:10/09:40 野々村和晃 Nakayama isomorphisms for the maximal quotient ring of a left Harada ring 09:50/10:20 本瀬香（弘前大理工）

Gaussian sums are just characters of the multiplicative group of a finite field

10:30/11:20 Rickard, Jeremy (Univ. of Bristol)

Derived Categories in Representation Theory, II

11:30/12:00 巧刀刀直子（愛知教育大） Some topics on derived equivalent blocks of finite groups

12:10/12:40 宇佐美陽子（お茶の水大理） On the principal 3-blocks of the Chevalley groups G_{2}(q)

14:00/14:30 加藤希理子（大阪女子大理） Morphisms represented by monomorphisms

14:40/15:10 荒谷督司（岡山大理） Remarks on a transitivity of exceptional sequences

15:20/16:10 Van den Bergh, Michel (Limburgs Universitair Centrum) Non-commutative Del Pezzo surfaces, II

16:30/17:00 中島晴久（城西大理） Relative invariants of groups acting on Krull domains

17:10/17:40 吉野雄二（岡山大理） Frobenius and quasi-Frobenius property for mod \underline{C}

17:50/18:20 高橋亮（岡山大理） On the decomosablility of a syzygy of the residue field

９月５日（日）

09:10/09:40 越谷重夫（千葉大理） On indecomposability of a module given by Brauer construction

09:50/10:40 Rickard, Jeremy (Univ. of Bristol)

Derived Categories in Representation Theory, III

10:50/11:40 Van den Bergh, Michel (Limburgs Universitair Centrum) Non-commutative crepant resolutions

11:50/12:30 伊山修（兵庫県立大理） Higher dimensional Auslander-Reiten theory on maximal orthogonal subcategories